OpenStudy (anonymous):

PLZ HELP write b^2 * b^5 without exponents

OpenStudy (anonymous):

b*b*b*b*b*b*b*

OpenStudy (anonymous):

You two are so fast. :)

OpenStudy (anonymous):

@MaryLovell15 i don't understand

OpenStudy (anonymous):

b^2=b*b b^5=b*b*b*b*b

OpenStudy (anonymous):

ohhhh ok ok @MaryLovell15